OpenStudy (anonymous):

The tangent line to the vertex of p(x)=4x2+8x+1 is:

OpenStudy (turingtest):

the vertex of a parabola is always a max or min value. In this case because the parabola points upwards it is a min. At max/min of p(x) p'(x)=8x+8=0 x=-1 p(-1)=4-8+1=-3 for point-slope form: y+3=p'(-1)(x+1)=(0)(x+1)=0 y=-3

OpenStudy (anonymous):

the vertex occurs where dp/dx=0

OpenStudy (anonymous):

this is dp(x)/dx=0=8x+8 so, x=-1

OpenStudy (anonymous):

the corresponding y coordinate is p(-1)=4(-1)2+8(-1)+1=-3

OpenStudy (anonymous):

the vetx is (-1,-3)

OpenStudy (anonymous):

the tangent is a line with forumla y=mx+b. it passes through the point (-1,-3) and it's slope (m) is given by dp/dx at x=-1